OpenStudy (anonymous):

The cost C (in dollars) of manufacturing x wheels at Mai's Bicycle Supply is given by the function c(x)=0.5x^2-170x+23,870. . How many wheels must be made to minimize the cost? Do not round your answer

OpenStudy (er.mohd.amir):

use diff.calculas d/dx{c(x)=0.5*2*x-170 equating to zero give x=170

OpenStudy (anonymous):

@Er.Mohd.AMIR ty

OpenStudy (mathmate):



OpenStudy (mathmate):

If you do not use calculus, it can be found by completing the square: c(x)=ax^2-bx+c then the minimum occurs at b/(2a) = 170/(2*(1/2))=170.